OpenStudy (makuza123):

Solve the following system of equations. Express your answer as an ordered pair in the format (a,b), with no spaces between the numbers or symbols. 2x+7y=-1 4x-3y=-19

OpenStudy (texaschic101):

2x + 7y = -1......(- 2) 4x - 3y = - 19 ------------------- - 4x - 14y = 2 4x - 3y = - 19 -------------------- - 17y = - 17 y = 1 2x + 7y = - 1 2x + 7(1) = - 1 2x + 7 = - 1 2x = - 8 x = - 4 (- 4, 1)

OpenStudy (texaschic101):

any questions ?

OpenStudy (makuza123):

hold on.. how you got (-2)?

OpenStudy (texaschic101):

basically, I multiplied 1st eq by - 2....then adding the equations, thus, eliminating the x's and solving for y. Once I found y, I subbed that back into an original equation to find x.

OpenStudy (texaschic101):

understand ?

OpenStudy (texaschic101):

this is elimination method...in this case, easier then substitution...so we r not stuck dealing with fractions.

OpenStudy (makuza123):

alright, and thank you for your help

OpenStudy (texaschic101):

yw